### v2

Abstract
In this work, we explore the effect of neutrino nonstandard interactions (NSI) involving the charm quark at SND@LHC. Using an effective description of new physics in terms of four-fermion operators involving a charm quark, we constrain the Wilson coefficients of the effective interaction from two and three-body charmed meson decays. In our fit, we include charmed meson decays not only to pseudoscalar final states but also to vector final states and include decays to the $\eta$ and $\eta^{\prime}$ final states. We also consider constraints from charmed baryon decays. We then study the effect of new physics in neutrino scattering processes, involving charm production at SND@LHC, for various benchmark new physics couplings obtained from the low energy fits. Finally, we also study the effects of lepton universality violation (LUV) assuming that the new physics coupling is not lepton universal.
1 Introduction
The observation of neutrino masses and mixing indicate physics beyond the Standard Model (SM). Therefore, it is reasonable to hypothesize that neutrinos may have new interactions beyond the SM. These interactions, known as neutrino nonstandard interactions (NSI), can be explored in specific models or in a model-independent framework in terms of four-fermion operators. The interactions may be purely leptonic or semileptonic where in the latter case a quark current and a leptonic current are involved in the effective interaction. In typical neutrino experiments, NSI involving the first-generation quarks are involved. In many models of BSM physics, the new physics (NP) effects are more pronounced in the heavier generations. NSI involving the heavy quarks and leptons are particularly interesting in these models. Also, given that there are hints of NP in decays of heavy quarks, a program to explore NSI involving heavy quarks is quite compelling. New high-energy neutrino scattering experiments that will produce charm quarks offer a unique opportunity to study NSI with heavy quarks.
The SM is lepton-flavor universal, i.e. the electroweak gauge interactions of the SM apply identically to all three flavors of leptons. Violation of this universality, known as lepton universality violation (LUV), is a crucial test of the SM. Evidence of LUV in $B$-mesons decays has generated significant interest in testing for LUV in various decays. In $B$ decays, hints of LUV have been observed in the charged-current quark-level transition $b\to c\ell^{-}\bar{\nu}$, where $\ell$ can be an $e,\mu,$ or $\tau$. At the mesonic level, since the hadronic part of such a decay rate is plagued by uncertainties from form factors and the non-perturbative nature of QCD, one considers ratios of decay rates as relatively cleaner tests for LUV.
[…]
Current experimental measurements also allow a 3-5% LUV in the ratio $R_{D^{(*)}}\equiv{\cal B}({\bar{B}}\to D^{(*)}\mu^{-}{\bar{\nu}}_{\mu})/{\cal B}({\bar{B}}\to D^{(*)}e^{-}{\bar{\nu}}_{e})$, which is expected to be $\sim 1$ in the SM. For the interested reader, we summarize the SM predictions for these observables and their experimental measurements in Appendix A.
Hints of NP are not restricted to charged-current $B$ decays. A recent first measurement of the branching ratio ${\cal B}(B^{+}\to K^{+}\nu\bar{\nu})=(2.3\pm 0.7)\times 10^{-5}$ by the Belle II experiment Belle-II:2023esi is $2.7\sigma$ higher than the SM expectation ${\cal B}(B^{+}\to K^{+}\nu\bar{\nu})_{\rm SM}=(5.58\pm 0.38)\times 10^{-6}$ Parrott:2022zte . Furthermore, even though the recently updated measurements of $R_{K^{(*)}}={\cal B}(B\to K^{(*)}\mu^{+}\mu^{-})/{\cal B}(B\to K^{(*)}e^{+}e^{-})$ are now fully consistent with their SM expectations LHCb:2022qnv , the individual branching fractions in both the electron and muon channels remain discrepant Capdevila:2023yhq ). Joint explanations of all these anomalies, both in model-dependent as well as model-independent approaches, favor NP that affects the heavier generations of quarks and leptons (see for example Refs. Bhattacharya:2016mcc ; Bhattacharya:2014wla .)
While many of these anomalies have been observed in low-energy data, if NP is involved, it should also affect quark-lepton interactions at higher energy scales. In this paper, we study the effects of NP, specifically in the neutrino-quark scattering process $\nu_{\ell}+q\to q^{\prime}+\ell^{-}$ where $\ell$ can be $e,\mu,$ or $\tau$. The effects of LUV NP in the scattering of $\nu_{\tau}$ off light quarks were studied in Refs. Rashed:2012bd ; Rashed:2013dba ; Liu:2015rqa . The parameter space of NP couplings and energy scales for $\nu_{\tau}$ scattering off light quarks can be constrained using data from hadronic tau decays. However, since the light leptons do not decay hadronically, similar constraints do not appear for $\nu_{e}$ and $\nu_{\mu}$. Instead, if the scattering of $\nu_{\ell}$ off of a light quark ($q$) produces a heavy quark ($h$) then the corresponding parameter space can be constrained by studying semileptonic transitions of a heavy quark to a light quark, $h\to q\ell^{-}{\bar{\nu}}$.
[…]
Neutrinos reaching the SND@LHC detector are energetic enough to produce heavy quarks, such as the charm quark, in the final state. This allows SND@LHC to be a laboratory for testing LUV effects in quark-neutrino scattering processes involving a heavy quark. The SND@LHC detector also aims at exploring the possibility of detecting new particles that scatter similar to neutrinos, such as light dark matter (LDM) particles which interact with SM particles through portal mediators. This study focuses on estimating the sensitivity of the SND@LHC detector in detecting new-physics phenomena, considering various benchmark new-physics couplings derived from low-energy fits.
The SND@LHC detector comprises a target region followed by the muon system. The detector pseudorapidity range spans from 7.2 to 8.4. The target consists of five walls of emulsion cloud chambers (ECC) followed by planes of Scintillating Fiber (SciFi) trackers. Each wall comprises 60 emulsion films interleaved with 59 tungsten plates, each 1 mm thick, serving as the target material. The ECC provides micrometric accuracy for measuring charged-particle tracks and reconstructing vertices of neutrino interactions. The reconstruction of particles and showers spanning several emulsion bricks is aided by the interleaved layers of SciFi, that provide accurate time stamps. When combined with the ECC walls used as radiators, the SciFi detector also serves as a sampling calorimeter to measure the energy of electromagnetic showers. Hadronic showers start developing in the target volume, but then they are fully contained by a hadronic calorimeter composed of alternating layers of 20 cm thick iron walls and 1 cm thick scintillating bars.
A key feature of SND@LHC is its high efficiency in identifying neutrino flavors. Electron neutrinos can be identified via their electromagnetic showers with 99% efficiency. Charged-current muon neutrino interactions can be detected with 69% efficiency by requiring the presence of at least one muon track. Neutral current interactions are correctly tagged with 99% efficiency. Thanks to the micrometric resolution of the ECC, tau neutrinos can be identified via a decay vertex that is displaced from the primary interaction point. The $\nu_{\tau}$ detection efficiency for SND@LHC ranges from 48% to 54% depending on the decay mode of the $\tau$ lepton Ahdida:2750060 ; SNDLHC:2022ihg .
[…]
The paper is organized in the following manner. In section II, we describe the formalism of the effective Hamiltonian and discuss several UV complete models from which the effective operators in the Hamiltonian can emerge. In section III, we discuss the low-energy flavor observables, perform fits to constrain the WCs, and choose benchmark values for the neutrino scattering study. In section IV, we consider the neutrino scattering and discuss the level of significance at which LUV effects can be detected at SND@LHC. We summarize and present our conclusions in section V.
2 Effective Hamiltonian
We begin by considering an effective Hamiltonian with dimension 6 four-fermion operators and Wilson coefficients that encode high-energy NP. Using the effective Hamiltonian we choose benchmark values for Wilson coefficients using constraints from low-energy observables including leptonic and semileptonic decays of $D$ mesons. The benchmark values are then used to study the effects of NP on neutrino scattering. The effective Hamiltonian that contributes to the scattering process $\nu_{\ell}+q\to c+\ell^{-}$ can be expressed as Falkowski:2021bkq ; Kopp:2024yvh :
$H_{\rm eff}$ $=$ $\frac{4G_{F}V_{cq}}{\sqrt{2}}\left\{(1+\epsilon_{L})_{\alpha\beta}[{\bar{c}}\gamma_{\mu}P_{L}q][{{\bar{\ell}}}_{\alpha}\gamma^{\mu}P_{L}\nu_{\beta}]+(\epsilon_{R})_{\alpha\beta}[{\bar{c}}\gamma_{\mu}P_{R}q][{{\bar{\ell}}}_{\alpha}\gamma^{\mu}P_{L}\nu_{\beta}]\right.~{}$ (1)
$\hskip 28.45274pt+~{}(\epsilon_{S})_{\alpha\beta}[{\bar{c}}q][{{\bar{\ell}}}_{\alpha}P_{L}\nu_{\beta}]+(\epsilon_{P})_{\alpha\beta}[{\bar{c}}\gamma^{5}q][{{\bar{\ell}}}_{\alpha}P_{L}\nu_{\beta}]~{}$
$\hskip 142.26378pt\left.+~{}(\epsilon_{T})_{\alpha\beta}[{\bar{c}}\sigma_{\mu\nu}P_{L}q][{\bar{\ell}}_{\alpha}\sigma^{\mu\nu}P_{L}\nu_{\beta}]\right\}\,+~{}{\rm h.c.},~{}~{}$
where $\alpha,\beta$ are lepton flavor indices representing the three flavors $e,\mu$, and $\tau$, and $q=s,d$. Here $P_{R}$ and $P_{L}$ are respectively the right and left projection operators $(1\pm\gamma^{5})/2$, $G_{F}$ refers to the Fermi constant, $V_{cq}$ refers to the appropriate CKM matrix element, and $(\epsilon_{X})_{\alpha\beta}$ refer to the NP Wilson coefficients where $X$ can be $L,R,S,P$, or $T$ for left-handed vector, right-handed vector, scalar, pseudoscalar, or tensor. Here the neutrinos are left-handed Dirac particles and $X$ refers to the Lorentz structure of the quark current.
To obtain benchmarks for $(\epsilon_{X})_{\alpha\beta}$ we will study processes where the quark-level transition is of the type $({\bar{c}}q)({\bar{\ell}}\nu)$, specifically two- and three-body decays of $D$ and $D^{*}$ mesons. Now, since neutrinos are not directly detected at collider experiments the final state neutrino flavor in a two- or three-body meson decay is unknown. Any effect of neutrino flavor gets washed out in summing over the final-state neutrino flavor. While in principle NSI can lead to lepton flavor violation when $(\epsilon_{X})_{\alpha\beta}$ is nonzero for $\alpha\neq\beta$, here we will always sum over the neutrino flavor index $\beta$. The resulting Hamiltonian is effectively flavor diagonal and can be expressed as
$H_{\rm eff}$ $=$ $\frac{4G_{F}V_{cq}}{\sqrt{2}}\left\{(1+\epsilon_{L})[{\bar{c}}\gamma_{\mu}P_{L}q][{{\bar{\ell}}}\gamma^{\mu}P_{L}\nu]+\epsilon_{R}[{\bar{c}}\gamma_{\mu}P_{R}q][{{\bar{\ell}}}\gamma^{\mu}P_{L}\nu]\right.~{}$ (2)
$\hskip 5.69054pt\left.+~{}\epsilon_{S}[{\bar{c}}q][{{\bar{\ell}}}_{\alpha}P_{L}\nu_{\beta}]+\epsilon_{P}[{\bar{c}}\gamma^{5}q][{\bar{\ell}}P_{L}\nu]+\epsilon_{T}[{\bar{c}}\sigma_{\mu\nu}P_{L}q][{\bar{\ell}}\sigma^{\mu\nu}P_{L}\nu]\right\}\,+~{}{\rm h.c.},~{}~{}$
where without the loss of generality here and in what follows we have suppressed the charge-lepton flavor index $\alpha$ on the effective WCs $\epsilon_{X}$ ($X=L,R,S,P,T$) obtained after summing over the neutrino flavor index $\beta$,
$\epsilon_{X}^{\alpha}~{}=~{}\sum\limits_{\beta}^{e,\mu,\tau}(\epsilon_{X})_{\alpha\beta}~{}.~{}~{}$ (3)
It is common to use a different nomenclature and basis of WCs, especially for the scalar and pseudoscalar operators which can instead be written in terms of left- and right-handed operators. This is done by rewriting the above Hamiltonian as follows
$H_{\rm eff}$ $=$ $\frac{4G_{F}V_{cq}}{\sqrt{2}}\left\{(1+V_{L})[{\bar{c}}\gamma_{\mu}P_{L}q][{{\bar{\ell}}}\gamma^{\mu}P_{L}\nu]+V_{R}[{\bar{c}}\gamma_{\mu}P_{R}q][{{\bar{\ell}}}\gamma^{\mu}P_{L}\nu]\right.~{}$ (4)
$\hskip 5.69054pt\left.+~{}S_{L}[{\bar{c}}P_{L}q][{{\bar{\ell}}}P_{L}\nu]+S_{R}[{\bar{c}}P_{R}q][{\bar{\ell}}P_{L}\nu]+T_{L}[{\bar{c}}\sigma_{\mu\nu}P_{L}q][{\bar{\ell}}\sigma^{\mu\nu}P_{L}\nu]\right\}\,+~{}{\rm h.c.},~{}~{}$
where $X_{Y}$ represent the effective NP WCs with $X=S,V,T$ referring to scalar, vector, or tensor and $Y=L,R$ referring to left-handed and right-handed quark currents. We have once again suppressed the charged-lepton flavor index, $\alpha$, in Eq. (4). By comparing Eqs. (2) and (4) one can relate the two sets of WCs as follows,
$V_{L(R)}~{}=~{}\epsilon_{L(R)}~{},~{}~{}S_{L}~{}=~{}\epsilon_{S}-\epsilon_{P}~{},~{}~{}S_{R}~{}=~{}\epsilon_{S}+\epsilon_{P}~{}.~{}~{}T_{L}~{}=~{}\epsilon_{T}~{},~{}~{}$ (5)
Thus, constraints from meson decay experiments can be directly translated into bounds on five WCs, $V_{L},V_{R},S_{L},S_{R}$, and $T_{L}$ for each lepton flavor $e,\mu$, and $\tau$.
The effective Hamiltonian given in Eq. (4) above can come from various ultraviolet-complete models. We discuss two specific examples – the leptoquark model and the vector boson ($W^{\prime}$) model. Here we demonstrate how the Wilson coefficients in Eq. (4) can be expressed in terms of tree-level couplings within these ultraviolet-complete models.
The interaction of a singlet leptoquark, $S_{1}(\bar{3},1,1/3)$, with SM fermions can be expressed as Buchmuller:1986zs ,
${\cal L}^{\rm LQ}=\left(g_{1L}^{ij}\,\bar{Q}_{iL}^{c}i\sigma_{2}L_{jL}+g_{1R}^{ij}\,\bar{u}_{iR}^{c}\ell_{jR}\right)S_{1}\,+~{}{\rm h.c.},$ (6)
where $Q_{i}$ and $L_{j}$ represent the left-handed quark and lepton doublets, $u_{iR}$ and $d_{iR}$ are the right-handed up-type and down-type quark singlets and $\ell_{jR}$ represents the right-handed charged lepton singlets. Indices $i$ and $j$ denote the generations of quarks and leptons. Integrating out the heavy $S_{1}$ leptoquark leads to,
${\cal L}_{\rm eff}$ $=$ $-\frac{g^{ij}_{1L}g^{kl*}_{1R}}{M^{2}_{S_{1}}}\left(\bar{Q}_{iL}^{c}i\sigma_{2}L_{jL}\right)\left({{\bar{\ell}}}_{lR}u^{c}_{kR}\right)\,+~{}{\rm h.c.},~{}~{}$ (7)
$=$ $-\frac{g^{ij}_{1L}g^{kl*}_{1R}}{M^{2}_{S_{1}}}\left\{\left({\bar{u}}^{c}_{iL}\ell_{jL}\right)\left({{\bar{\ell}}}_{lR}u^{c}_{kR}\right)-\left({\bar{d}}^{c}_{iL}\nu_{jL}\right)\left({{\bar{\ell}}}_{lR}u^{c}_{kR}\right)\right\}\,+~{}{\rm h.c.}~{}~{}$ (8)
The second of the two terms above, after Fierz transformation, gives rise to the following effective Hamiltonian,
${\cal H}_{\rm eff}$ $=$ $\frac{g^{2l}_{1L}g^{2l*}_{1R}}{2M^{2}_{S_{1}}}\left[\left({\bar{c}}P_{L}s\right)\left({{\bar{\ell}}}P_{L}\nu\right)+\frac{1}{4}\left({\bar{c}}\sigma^{\mu\nu}P_{L}s\right)\left({{\bar{\ell}}}\sigma_{\mu\nu}P_{L}\nu\right)\right]\,+~{}{\rm h.c.}~{}~{}$ (9)
Comparing this with the effective Hamiltonian of Eq. (4) and restricting ourselves to second-generation quarks, we can express the coefficients $S_{L}$ and $T_{L}$ as follows,
$S_{L}=\frac{1}{2\sqrt{2}G_{F}V_{cs}}\frac{g_{1L}^{2l}g_{1R}^{2l*}}{2M_{S_{1}}^{2}},\qquad T_{L}=\frac{1}{2\sqrt{2}G_{F}V_{cs}}\frac{g_{1L}^{2l}g_{1R}^{2l*}}{8M_{S_{1}}^{2}}.\,$ (10)
Let us now consider the following Lagrangian, encoding the interaction of a $W^{\prime}$ boson with leptons and quarks:
${\cal L}^{W^{\prime}}$ $=$ $\frac{g}{\sqrt{2}}\left[V_{ij}\left\{\bar{u}_{i}\gamma^{\mu}(g^{ij}_{L}P_{L}+g^{ij}_{R}P_{R})d_{j}\right\}+g^{\nu_{\ell}\ell}\left({\bar{\ell}}\gamma^{\mu}P_{L}\nu_{\ell}\right)\right]W^{\prime+}_{\mu}\,+~{}{\rm h.c.},~{}~{}$ (11)
where $g$ is the SM Weak coupling, $V_{ij}$ represents the relevant CKM matrix element, $g^{ij}_{X}$ ($X=L,R$) and $g^{\nu_{\ell}\ell}$ are the NP couplings of the $W^{\prime}$ boson with the SM quarks and leptons. Using the definition of the fermi constant, $G_{F}/\sqrt{2}=g^{2}/8m^{2}_{W}$, and integrating out the $W^{\prime}$ leads to the following effective Hamiltonian.
${\cal H}_{\rm eff}$ $=$ $\frac{4G_{F}V_{cq}}{\sqrt{2}}\left[\bar{c}_{i}\gamma^{\mu}\left(\frac{M_{W}^{2}}{M_{W^{\prime}}^{2}}g^{cq}_{L}P_{L}+\frac{M_{W}^{2}}{M_{W^{\prime}}^{2}}g^{cq}_{R}P_{R}\right)q\right]\left[g^{\nu_{\ell}\ell}{\bar{\ell}}\gamma_{\mu}P_{L}\nu_{\ell}\right]\,+~{}{\rm h.c.}~{}~{}$ (12)
Comparing Eq. (12) with Eq. (4) we obtain the following relations:
$V_{L}~{}=~{}\frac{M_{W}^{2}}{M_{W^{\prime}}^{2}}g^{cq}_{L}\sum\limits_{\nu_{\ell}}g^{\nu_{\ell}\ell},~{}~{}~{}~{}~{}V_{R}~{}=~{}\frac{M_{W}^{2}}{M_{W^{\prime}}^{2}}g^{cq}_{R}\sum\limits_{\nu_{\ell}}g^{\nu_{\ell}\ell}.$ (13)
3 Flavor Observables
In this section, we discuss the low-energy fits to the WCs of the effective Lagrangian from low-energy data and select a few benchmarks for NP. There are two main types of observables, listed below, that we are interested in for benchmarking.
  • Two-body decays: $D_{s}^{+}\to\ell^{+}\nu,~{}D^{+}\to\ell^{+}\nu$
  • Three-body decays: $D\to K\ell\nu$$D\to\pi\ell\nu$,  $D\to K^{*}\ell\nu$$D\to\rho\ell^{+}\nu$,  $D_{s}^{+}\to\phi\ell^{+}\nu$,  $D_{s}^{+}\to\eta^{(\prime)\ell\nu}$$\Lambda_{c}\to\Lambda\ell\nu$
In the following sections, we will discuss benchmarks obtained from the above list.
3.1 Two-body meson decays
The decay rate for the process $M^{+}_{ij}\to\ell^{+}\nu$, where $M^{+}_{ij}$ represents a pseudoscalar $q^{i}_{u}{\bar{q}_{d}}^{j}$ meson, can be expressed in the basis of Eq. (4) as:
$\Gamma_{M^{+}_{ij}\to\ell^{+}\nu}~{}=~{}\frac{G^{2}_{F}m_{M}m^{2}_{\ell}}{8\pi}|V_{ij}|^{2}f^{2}_{M}\left(1-\frac{m^{2}_{\ell}}{m^{2}_{M}}\right)^{2}\left\{|1+V_{L}-V_{R}|^{2}+\frac{|S_{L}-S_{R}|^{2}m^{4}_{M}}{m^{2}_{\ell}(m_{i}+m_{j})^{2}}\right\},~{}~{}$ (14)
where $m_{i}$ is the mass of the $q^{i}_{u,d}$, $V_{ij}$ represents the relevant CKM matrix element, and other symbols carry their usual meaning. All numerical inputs are listed in Table 7 in Appendix C. The SM predictions using these parameters are given in Table 1.
Similarly, the decay rate for the process $M^{*+}\to\ell^{+}\nu$, where $M^{*+}$ represents a vector meson, can be expressed as:
$\Gamma_{M^{*+}\to\ell^{+}\nu}~{}=~{}\frac{G^{2}_{F}|V_{q_{u}q_{d}}|^{2}f_{M}^{2}m^{3}_{M}}{192\pi}\left(1-\frac{m^{2}_{\ell}}{m^{2}_{M}}\right)^{2}\left(2+\frac{m^{2}_{\ell}}{m^{2}_{M}}\right)^{2}|1+V_{L}+V_{R}|^{2}.~{}~{}$ (15)
The above result ignores a tensor decay constant that a vector meson may have. Note that, since the SM does not have a tensor current, the corresponding decay constant for a vector meson can not be measured in a model-independent fashion. Hence we remove this contribution for brevity.
3.2 Three-body meson decays
The decay rate of the semileptonic three body meson decays $M_{i}\to M_{j}\ell\nu$ in the basis of the Hamiltonian in Eq. (4) can be written as:
$\frac{d\Gamma}{dq^{2}}$ $=$ $\frac{G_{F}^{2}|V_{ij}|^{2}}{192\pi^{3}m_{M_{i}}^{3}}q^{2}\sqrt{\lambda(m_{M_{i}}^{2},m_{M_{j}}^{2},q^{2})}\left(1-\frac{m_{\ell}^{2}}{q^{2}}\right)^{2}\times$ (16)
$\bigg{\{}|1+V_{L}+V_{R}|^{2}\left[\left(1+\frac{m_{\ell}^{2}}{2q^{2}}\right)h_{V,0}^{2}(q^{2})+~{}\frac{3}{2}\frac{m_{\ell}^{2}}{q^{2}}h_{V,t}^{~{}2}(q^{2})\right]+\frac{3}{2}|S_{L}+S_{R}|^{2}h_{S}^{2}(q^{2})$
$+~{}3\mathcal{R}\rm{e}\left[(1+V_{L}+V_{R})(S_{L}^{*}+S_{R}^{*})\right]\frac{m_{\ell}}{\sqrt{q^{2}}}h_{S}(q^{2})h_{V,t}(q^{2})\bigg{\}},$
where $\lambda(x,y,z)=x^{2}+y^{2}+z^{2}-2(xy+yz+zx)$.
The semileptonic three-body decay rate for $M_{i}\to M_{j}^{*}\ell\nu$ decays is expressed as:
$\frac{d\Gamma}{dq^{2}}$ $=$ $\frac{G_{F}^{2}|V_{ij}|^{2}}{192\pi^{3}m_{M_{i}}^{3}}q^{2}\sqrt{\lambda(m_{M_{i}}^{2},m_{M_{j}^{*}}^{2},q^{2})}\left(1-\frac{m_{\ell}^{2}}{q^{2}}\right)^{2}\times$ (17)
$\bigg{\{}(|1+V_{L}|^{2}+|V_{R}|^{2})\left(\left(1+\frac{m_{\ell}^{2}}{2q^{2}}\right)(H_{V,+}^{2}(q^{2})+H_{V,-}^{2}(q^{2})+H_{V,0}^{2}(q^{2}))+\frac{3}{2}\frac{m_{\ell}^{2}}{q^{2}}H_{V,t}^{2}\right)$
$-2\mathcal{R}e\left[(1+V_{L})V_{R}^{*}\right]\left(\left(1+\frac{m_{\ell}^{2}}{2q^{2}}\right)(2H_{V,+}^{2}(q^{2})H_{V,-}^{2}(q^{2})+H_{V,0}^{2}(q^{2}))+\frac{3}{2}\frac{m_{\ell}^{2}}{q^{2}}H_{V,t}^{2}(q^{2})\right)$
$+\frac{3}{2}|S_{L}-S_{R}|^{2}H_{S}^{2}(q^{2})+3\mathcal{R}e\left[(1+V_{L}-V_{R})(S_{L}^{*}-S_{R}^{*})\right]\frac{m_{\ell}}{\sqrt{q^{2}}}H_{S}(q^{2})H_{V,t}(q^{2})\bigg{\}}.$
The above hadronic amplitudes in terms of the form factors are provided in Appendix B and the form factor parameters and other numerical inputs are detailed in Appendix C.
3.3 Three-body baryonic decays
In this work we also consider the semileptonic decay of charmed baryon $\Lambda_{c}\to\Lambda\ell\nu$. The decay rate in terms of new physics Wilson coefficients as defined in Eq. (4) has been adopted from Ref. Shivashankara:2015cta ; Datta:2017aue for the $\Lambda_{b}\to\Lambda_{c}\ell\nu$ decay as :
$\frac{d\Gamma}{dq^{2}}=\frac{G_{F}^{2}|V_{cs}|^{2}q^{2}}{384\pi^{3}m_{\Lambda_{c}}^{3}}\sqrt{Q_{+}Q_{-}}\left(1-\frac{m_{\ell}^{2}}{q^{2}}\right)^{2}\left(A_{1}^{VA}+\frac{m_{\ell}^{2}}{2q^{2}}A_{2}^{VA}+\frac{3}{2}A_{3}^{SP}+\frac{3m_{\ell}}{\sqrt{q^{2}}}A_{5}^{VA-SP}\right),$ (18)
where $Q_{\pm}=(m_{\Lambda_{c}}\pm m_{\Lambda})^{2}-q^{2}$ and the amplitudes $A_{i}^{X}$ in terms of the form factors and WCs can be found in Ref. Datta:2017aue .
Decay $\ell=e$ $\ell=\mu$ $\ell=\tau$
$\mathcal{B}^{\rm{SM}}$ (%) $\mathcal{B}^{\rm{meas}}$ (%) $\mathcal{B}^{\rm{SM}}$ (%) $\mathcal{B}^{\rm{meas}}$ (%) $\mathcal{B}^{\rm{SM}}$ (%) $\mathcal{B}^{\rm{meas}}$ (%)
$D^{+}_{s}\to\ell^{+}\nu_{\ell}$ $(1.28\pm 0.05)\times 10^{-5}$ $<8.3\times 10^{-3}$ $0.546\pm 0.020$ $0.543\pm 0.015$ $5.32\pm 0.20$ $5.32\pm 0.11$
$D^{+}\to\ell^{+}\nu_{\ell}$ $(8.85\pm 0.47)\times 10^{-7}$ $<8.8\times 10^{-4}$ $0.0376\pm 0.0020$ $0.0374\pm 0.0017$ $0.100\pm 0.005$ $0.120\pm 0.027$
$D^{+}\to\overline{K}^{0}\ell^{+}\nu_{\ell}$ $8.99\pm 0.37$ $8.72\pm 0.09$ $8.77\pm 0.36$ $8.76\pm 0.19$
$D^{0}\to K^{-}\ell^{+}\nu_{\ell}$ $3.54\pm 0.15$ $3.549\pm 0.026$ $3.45\pm 0.14$ $3.41\pm 0.04$
$D^{+}\to\pi^{0}\ell^{+}\nu_{\ell}$ $0.320\pm 0.031$ $0.372\pm 0.017$ $0.316\pm 0.030$ $0.350\pm 0.015$
$D^{0}\to\pi^{-}\ell^{+}\nu_{\ell}$ $0.249\pm 0.024$ $0.291\pm 0.004$ $0.245\pm 0.023$ $0.267\pm 0.012$
$D^{+}\to\overline{K}^{*}(892)^{0}\ell^{+}\nu_{\ell}$ $5.84\pm 1.61$ $5.40\pm 0.10$ $5.52\pm 1.52$ $5.27\pm 0.15$
$D^{0}\to K^{*}(892)^{-}\ell^{+}\nu_{\ell}$ $2.30\pm 0.64$ $2.15\pm 0.16$ $2.18\pm 0.60$ $1.89\pm 0.24$
$D^{+}\to\rho^{0}\ell^{+}\nu_{\ell}$ $0.212\pm 0.048$ $0.19\pm 0.01$ $0.203\pm 0.046$ $0.24\pm 0.04$
$D^{0}\to\rho^{-}\ell^{+}\nu_{\ell}$ $0.171\pm 0.039$ $0.150\pm 0.012$ $0.163\pm 0.037$ $0.135\pm 0.013$
$D^{+}_{s}\to\phi\ell^{+}\nu_{\ell}$ $2.68\pm 0.26$ $2.39\pm 0.16$ $2.53\pm 0.23$ $1.9\pm 0.5$
$D^{+}_{s}\to\eta\ell^{+}\nu_{\ell}$ $2.96\pm 0.45$ $2.26\pm 0.05$ $2.91\pm 0.45$ $2.4\pm 0.5$
$D^{+}_{s}\to\eta^{\prime}(958)\ell^{+}\nu_{\ell}$ $0.909\pm 0.41$ $0.8\pm 0.04$ $0.869\pm 0.41$ $1.1\pm 0.5$
$\Lambda_{c}^{+}\to\Lambda\ell^{+}\nu_{\ell}$ $3.92\pm 0.63$ $3.56\pm 0.13$ $3.80\pm 0.60$ $3.48\pm 0.17$
Table 1: List of observables used to constrain the new physics coefficients along with their SM predictions and measured branching fractions. The measured values are taken from the PDG ParticleDataGroup:2022pth while the SM predictions are computed using the form factors defined in Appendix C and the CKM matrix elements listed in Table 7.
3.4 Low energy fit results
We fit the new physics Wilson coefficients in Eq. (4) to the observables listed in Table 1 separately for the electron and muon modes. In Table 1, we list the SM expectation and the measured values of the branching fractions for each decay mode. For our numerical analysis, we take the values of $D$-meson decay constants and lattice determinations of relevant CKM matrix elements from the FLAG Review FlavourLatticeAveragingGroupFLAG:2021npn  [^1] Ref. Bolognani:2024cmr points out that the inclusion of a universal electroweak correction in the Wilson coefficient $\epsilon_{L}$ leads to a $\sim$ 1% increase in $\epsilon_{L}$. This also leads to a smaller value of $|V_{cs}|$ compared to the one obtained by the PDG ParticleDataGroup:2022pth . In this work, however, we use $\epsilon_{L}=1$ and $|V_{cs}|$ as given in the FLAG Review FlavourLatticeAveragingGroupFLAG:2021npn .. In Appendix C we list the relevant numerical inputs and form factors used in this article. Note that there are 12 measurements for the electron modes and 14 for the muon modes which are utilized to fit to the respective flavour specific Wilson coefficients. For the tau mode, however, we only have measurements for the two body leptonic decay of the $D_{(s)}$ meson. Furthermore, since the $\nu_{\tau}$ detection efficiency at SND is low compared to the other lepton flavour, it is challenging to obtain any meaningful constraint on the new physics parameters. Hence, we do not fit to tau measurements in our analysis.
The allowed parameter values for a specific model X corresponding to a specific lepton flavour is determined using
$\chi^{2}(X)=\sum_{i=1}^{n}\frac{(\mathcal{B}_{i}^{\rm{th}}(X)-\mathcal{B}_{i}^{\rm{meas}})^{2}}{\sigma_{i}^{2}},$ (19)
where n is the total number of observables, $\mathcal{B}^{\rm{th}}$ is the theoretical branching fraction, $\mathcal{B}^{\rm{meas}}$ is the corresponding experimental measurement and $\sigma$ is the uncertainty from measurement and theory added in quadrature. For a given model, the set of Wilson coefficients that minimize the above $\chi^{2}$ function are found with their respective $1\sigma$ uncertainties using the algorithm supplied by the MINUIT package James:1975dr ; iminuit as shown in Tables 23. We fit to both real and imaginary components of the Wilson coefficients denoted as $\epsilon_{X}={\rm Re}[\epsilon_{X}]+i\,{\rm Im}[\epsilon_{X}]$ where $X$ can be $L,R,S,P$. In each case, we define the pull with respect to the SM as $\sqrt{\chi^{2}(SM)-\chi^{2}(X)}$.
Model Fit Parameters $\chi^{2}$/dof ${\rm Re}\left[\epsilon_{L}\right]$ ${\rm Im}\left[\epsilon_{L}\right]$ ${\rm Re}\left[\epsilon_{R}\right]$ ${\rm Im}\left[\epsilon_{R}\right]$ ${\rm Re}\left[\epsilon_{S}\right]$ ${\rm Im}\left[\epsilon_{S}\right]$ ${\rm Re}\left[\epsilon_{P}\right]$ ${\rm Im}\left[\epsilon_{P}\right]$ pull
1 ${\rm Re}\left[\epsilon_{L}\right],{\rm Im}\left[\epsilon_{L}\right],{\rm Re}\left[\epsilon_{R}\right],{\rm Im}\left[\epsilon_{R}\right],$ 1.18/4 $-1.14(13)$ $-0.54(14)$ $-0.56(13)$ $0.32(14)$ $-0.53(14)$ $-0.17(40)$ $0.0(1.2)$ $0.0(1.2)$ 2.95
${\rm Re}\left[\epsilon_{S}\right],{\rm Im}\left[\epsilon_{S}\right],{\rm Re}\left[\epsilon_{P}\right],{\rm Im}\left[\epsilon_{P}\right]$
2 ${\rm Re}\left[\epsilon_{L}\right],{\rm Re}\left[\epsilon_{R}\right],{\rm Re}\left[\epsilon_{S}\right],{\rm Re}\left[\epsilon_{P}\right]$ 1.36/8 $-0.877(60)$ $-0.83(7)$ $0.58(10)$ $0.0(1.2)$ 2.92
3 ${\rm Re}\left[\epsilon_{L}\right],{\rm Im}\left[\epsilon_{L}\right]$ 9.8/10 $-0.0043(35)$ $0.014(25)$ 0.32
4 ${\rm Re}[\epsilon_{L}]$ 9.8/11 $-0.0042(130)$ 0.32
5 ${\rm Re}[\epsilon_{R}],{\rm Im}[\epsilon_{R}]$ 9.7/10 $0.006(13)$ $0.00(13)$ 0.45
6 ${\rm Re}[\epsilon_{R}]$ 9.7/11 $0.006(13)$ 0.45
7 ${\rm Re}[\epsilon_{L}],{\rm Im}[\epsilon_{L}],{\rm Re}[\epsilon_{R}],{\rm Im}[\epsilon_{R}]$ 8.2/8 $-0.32(12)$ $-0.70(19)$ $0.021(90)$ $-0.21(26)$ 1.3
8 ${\rm Re}[\epsilon_{L}],{\rm Re}[\epsilon_{R}]$ 8.2/10 $-0.028(24)$ $0.03(2)$ 1.3
9 ${\rm Re}[\epsilon_{S}],{\rm Im}[\epsilon_{S}]$ 8.15/5 $-0.098(230)$ $0.0(9)$ 0.52
10 ${\rm Re}[\epsilon_{S}]$ 8.15/6 $-0.098(230)$ 0.52
11 ${\rm Re}[\epsilon_{P}],{\rm Im}[\epsilon_{P}]$ 1.79/4 $0.0(6)$ $0.0(6)$ 0.004
12 ${\rm Re}[\epsilon_{P}]$ 1.79/5 $0.0(6)$ 0.004
13 ${\rm Re}[\epsilon_{S}],{\rm Im}[\epsilon_{S}],{\rm Re}[\epsilon_{P}],{\rm Im}[\epsilon_{P}]$ 9.63/8 $-0.098(290)$ $0.0(8)$ $-0.002(600)$ $0.0(6)$ 0.52
14 ${\rm Re}[\epsilon_{S}],{\rm Re}[\epsilon_{P}]$ 9.63/10 $-0.098(290)$ $-0.002(600)$ 0.52
15 ${\rm Re}[\epsilon_{L}],{\rm Im}[\epsilon_{L}],{\rm Re}[\epsilon_{R}],{\rm Im}[\epsilon_{R}],$ 1.18/6 $-0.73(60)$ $-0.49(17)$ $0.46(60)$ $0.45(18)$ $0.56(34)$ $-0.032(180)$ 2.95
${\rm Re}[\epsilon_{S}],{\rm Im}[\epsilon_{S}]$
16 ${\rm Re}[\epsilon_{L}],{\rm Re}[\epsilon_{R}],{\rm Re}[\epsilon_{S}]$ 1.35/8 $-0.88(6)$ $-0.83(7)$ $-0.58(11)$ 2.92
17 ${\rm Re}[\epsilon_{L}],{\rm Im}[\epsilon_{L}],{\rm Re}[\epsilon_{R}],{\rm Im}[\epsilon_{R}],$ 8.2/6 $-0.54(14)$ $0.86(18)$ $0.014(130)$ $0.026(180)$ $0.0(4)$ $0.0(3)$ 1.3
${\rm Re}[\epsilon_{P}],{\rm Im}[\epsilon_{P}]$
18 ${\rm Re}[\epsilon_{L}],{\rm Re}[\epsilon_{R}],{\rm Re}[\epsilon_{P}]$ 8.2/9 $-0.028(23)$ $0.030(24)$ $0.0(4)$ 1.3
Table 2: Fit results for different combinations of WCs in decays to electrons. The SM chi-square for the full set of observables listed in Table 1 is $\chi^{2}_{SM}/dof=9.90/12$ while for those affecting the scalar coupling only, $\chi^{2}_{SM,S}/dof=8.43/7$, and correspondingly for the pseudoscalar coupling, $\chi^{2}_{SM,P}/dof=1.79/6$.
Note that, while all observables considered in this work are sensitive to vector-type new physics, not all decay rates are affected by the scalar coupling $\epsilon_{S}$ or pseudoscalar coupling $\epsilon_{P}$. The two body leptonic decay of pseudoscalar mesons $M_{i}\to\ell\nu$ are purely sensitive to $\epsilon_{P}$ while that for the vector mesons are sensitive to neither $\epsilon_{S}$ nor $\epsilon_{P}$. The three body semileptonic decays of the form $M_{i}\to M_{j}\ell\nu$ are affected by $\epsilon_{S}$ only while $M_{i}\to M_{j}^{*}\ell\nu$ rates are sensitive to $\epsilon_{P}$ only. The charmed baryonic decay $\Lambda_{c}\to\Lambda\ell\nu$, however, are influenced by both scalar and pseudoscalar new physics. Hence, while performing the fits to models involving either $\epsilon_{S}$ or $\epsilon_{P}$ only, we drop the decay modes that are redundant to the analysis for that particular coupling. The SM chi-squares are calculated accordingly in order to estimate the pull with respect to new physics. As mentioned in the caption of Table 2, the SM chi-square value per degree of freedom ($dof$) for the full set of 12 observables listed in Table 1 is calculated to be $\chi^{2}_{SM}/dof=9.9/12$ while, for the observables sensitive to $\epsilon_{S}$ only, $\chi^{2}_{SM,S}/dof=8.43/7$, and, for those affected by $\epsilon_{P}$, $\chi^{2}_{SM,P}/dof=1.79/6$. Similarly, for the muon fit, the respective SM chi-squares are found to be $\chi^{2}_{SM}/dof=4.94/14$, $\chi^{2}_{SM,S}/dof=2.5/7$ and $\chi^{2}_{SM,P}/dof=2.7/8$.
The pseudoscalar coupling $\epsilon_{P}$ is tightly constrained from the two body decays $D_{s}^{+}\to\ell^{+}\nu_{\ell}$ and $D^{+}\to\ell^{+}\nu_{\ell}$. However, the lack of measurement of these decay rates for the electron channel keeps $\epsilon_{P}$ practically unconstrained as is reflected in the fit results in Table 2. In case of the muon, $\epsilon_{P}$ is better constrained due to the availability of the measurements of the two body decays and we obtained tightly constrained central values with small error bars as shown in Table 3. For the scalar and vector couplings, there is an interplay between the semileptonic mesonic three body decays and the charmed baryon decays resulting in very small to even $\mathcal{O}(1)$ central values for the real and imaginary parts of the parameters.
Model Fit Parameters $\chi^{2}$/dof ${\rm Re}\left[\epsilon_{L}\right]$ ${\rm Im}\left[\epsilon_{L}\right]$ ${\rm Re}\left[\epsilon_{R}\right]$ ${\rm Im}\left[\epsilon_{R}\right]$ ${\rm Re}\left[\epsilon_{S}\right]$ ${\rm Im}\left[\epsilon_{S}\right]$ ${\rm Re}\left[\epsilon_{P}\right]$ ${\rm Im}\left[\epsilon_{P}\right]$ pull
1 ${\rm Re}\left[\epsilon_{L}\right],{\rm Im}\left[\epsilon_{L}\right],{\rm Re}\left[\epsilon_{R}\right],{\rm Im}\left[\epsilon_{R}\right],$ 2.08/6 $-1.71(10)$ $0.36(14)$ $-0.14(9)$ $-0.39(14)$ $-0.36(13)$ $-0.011(8)$ $0.007(8)$ $-0.009(7)$ 1.69
${\rm Re}\left[\epsilon_{S}\right],{\rm Im}\left[\epsilon_{S}\right],{\rm Re}\left[\epsilon_{P}\right],{\rm Im}\left[\epsilon_{P}\right]$
2 ${\rm Re}\left[\epsilon_{L}\right],{\rm Re}\left[\epsilon_{R}\right],{\rm Re}\left[\epsilon_{S}\right],{\rm Re}\left[\epsilon_{P}\right]$ 2.14/10 $-0.097(10)$ $-0.049(14)$ $0.35(16)$ $-0.011(9)$ 1.67
3 ${\rm Re}\left[\epsilon_{L}\right],{\rm Im}\left[\epsilon_{L}\right]$ 4.87/12 $-1.98(1.10)$ $-0.18(1.70)$ 0.26
4 ${\rm Re}[\epsilon_{L}]$ 4.87/13 $-2.00(18)$ 0.26
5 ${\rm Re}[\epsilon_{R}],{\rm Im}[\epsilon_{R}]$ 4.83/12 $0.0038(110)$ $0.0(2)$ 0.34
6 ${\rm Re}[\epsilon_{R}]$ 4.83/12 $0.0038(110)$ 0.34
7 ${\rm Re}[\epsilon_{L}],{\rm Im}[\epsilon_{L}],{\rm Re}[\epsilon_{R}],{\rm Im}[\epsilon_{R}]$ 4.70/10 $-1.91(28)$ $-0.41(15)$ $0.00(28)$ $0.00(15)$ 0.48
8 ${\rm Re}[\epsilon_{L}],{\rm Re}[\epsilon_{R}]$ 4.70/12 $-0.004(12)$ $0.005(12)$ 0.48
9 ${\rm Re}[\epsilon_{S}],{\rm Im}[\epsilon_{S}]$ 1.24/5 $-0.31(6)$ $0.00(34)$ 1.12
10 ${\rm Re}[\epsilon_{S}]$ 1.24/6 $-0.31(6)$ 1.12
11 ${\rm Re}[\epsilon_{P}],{\rm Im}[\epsilon_{P}]$ 2.70/6 $-0.00014(13)$ $0.00000(13)$ 0.01
12 ${\rm Re}[\epsilon_{P}]$ 2.70/7 $-0.00014(13)$ 0.01
13 ${\rm Re}[\epsilon_{S}],{\rm Im}[\epsilon_{S}],{\rm Re}[\epsilon_{P}],{\rm Im}[\epsilon_{P}]$ 3.68/10 $-0.31(16)$ $0.00(33)$ $-0.00013(13)$ $0.00000(13)$ 1.12
14 ${\rm Re}[\epsilon_{S}],{\rm Re}[\epsilon_{P}]$ 3.68/12 $-0.31(16)$ $-0.00013(13)$ 1.12
15 ${\rm Re}[\epsilon_{L}],{\rm Im}[\epsilon_{L}],{\rm Re}[\epsilon_{R}],{\rm Im}[\epsilon_{R}],$ 2.65/8 $-0.11(26)$ $0.09(15)$ $-0.057(80)$ $-0.19(27)$ $0.36(31)$ $-0.04(40)$ 1.51
${\rm Re}[\epsilon_{S}],{\rm Im}[\epsilon_{S}]$
16 ${\rm Re}[\epsilon_{L}],{\rm Re}[\epsilon_{R}],{\rm Re}[\epsilon_{S}]$ 2.65/10 $-0.082(40)$ $-0.075(40)$ $0.36(15)$ 1.51
17 ${\rm Re}[\epsilon_{L}],{\rm Im}[\epsilon_{L}],{\rm Re}[\epsilon_{R}],{\rm Im}[\epsilon_{R}],$ 3.76/8 $-0.51(10)$ $-0.83(32)$ $0.016(100)$ $-0.028(320)$ $-0.007(400)$ $0.011(22)$ 1.09
${\rm Re}[\epsilon_{P}],{\rm Im}[\epsilon_{P}]$
18 ${\rm Re}[\epsilon_{L}],{\rm Re}[\epsilon_{R}],{\rm Re}[\epsilon_{P}]$ 3.76/10 $-0.031(30)$ $0.033(30)$ $-0.013(6)$ 1.09
Table 3: Similar to Table 2, but for the muon observables with $\chi^{2}_{SM}/dof=4.94/14$, $\chi^{2}_{SM,S}/dof=2.5/7$ and $\chi^{2}_{SM,P}/dof=2.7/8$.
4 Deep Inelastic Neutrino-Nucleon Scattering
In this section, we explore the sensitivity of SND@LHC to NP, as encapsulated in the WCs defined in Eq. (4). Given the high energy of the incoming neutrinos, neutrino detection proceeds through charged-current deep-inelastic scattering (DIS):
$\nu_{\ell}+N\to\ell+X,$ (20)
where $N=p,n$ represents a nucleon, $\ell=e\,,\mu\,,\tau$ and $X$ denotes any hadron state. However, NP contributes specifically to charm production, as shown in Fig. 1:
$\nu_{\ell}+N\to\ell+X_{c},$ (21)
where $X_{c}$ denotes a possible charm hadron state. In this analysis, we evaluate the confidence level with which the WCs obtained by fitting low-energy observables, as summarized in Tables 2 and 3, can be probed at SND@LHC.
The sensitivity is determined through a $\chi^{2}$ analysis for each type of neutrino using the following definition:
$\chi^{2}\,=\,\sum_{i=1}^{n}\left(\frac{N_{i}^{\rm SM+NP}-N_{i}^{\rm SM}}{\sqrt{N_{i}^{\rm SM+NP}}}\right)^{2}$ (22)
where $N_{i}^{\rm SM}$ are the number of differential events predicted by the SM and $N_{i}^{\rm SM+NP}$ are total number of differential events, including contributions from NP, for the $i^{\rm th}$ bin in the energy distribution.
We also assess the sensitivity to NP in the ratio $R_{\mu,e}\,=\,N_{\mu}/N_{e}$ for WCs given in Table 2 and Table 3, considering scenarios where NP effects are exclusively present in either muons or electrons at a time:
$\chi^{2}=\sum_{i=1}^{n}\left(\frac{R_{\mu,e}^{\rm SM+NP}-R_{\mu,e}^{\rm SM}}{\sigma_{R_{\mu,e}^{\rm SM+NP}}}\right)^{2},$ (23)
where the statistical error in the ratio $\sigma_{R_{\mu,e}}$ is calculated as $\sigma_{R_{\mu,e}}^{2}\equiv R_{\mu,e}^{2}\left({N_{e}}^{-1}+{N_{\mu}}^{-1}\right)$.
For signal or background, the total number of events is calculated from the DIS cross-section, $\sigma_{\nu N}$, as follows:
$N_{i}^{\rm model}$ $=$ ×$\mathcal{L}\times\sigma(pp\to\nu X)\times A_{g}\times P_{\rm int}(\sigma_{\nu N}^{\rm model}),$ (24)
$=$ $\Phi_{i,\nu}\times P_{\rm int}(\sigma_{\nu N}^{\rm model})$ (25)
where $\mathcal{L}=290$ fb-1 is the $pp$ luminosity, $\sigma(pp\to\nu X)$ is the cross-section for neutrino production in $pp$ collisions, $A_{g}$ is the geometrical acceptance of SND@LHC, and the product of these factors constitutes the incoming neutrino spectra $\Phi_{i,\nu}$ in the $i^{\rm th}$ energy bin. $P_{\rm int}(\sigma_{\nu N}^{\rm model})$ represents the interaction probability of a neutrino with the detector, which is a function of the DIS cross-section. Here, model refers to background (SM) or signal (SM+NP).
We take and adapt the spectra of neutrinos produced at the ATLAS LHC interaction point and impinging onto the SND@LHC detector from Ref. Ahdida:2750060 . These were obtained using the Pythia8, DPMJET and FLUKA library to simulate the production of neutrinos in $pp$ collisions, and their propagation through the machine elements until the location of the SND@LHC target. Only neutrinos within the SND@LHC geometrical acceptance $A_{g}$ are retained. These spectra are scaled up to an expected luminosity of $\mathcal{L}=290~{}\rm{fb}^{-1}$ and shown in Fig. 2.
Figure 1: Feynman diagram illustrating charged-current deep-inelastic scattering. This diagram with reversed arrows also applies to anti-neutrino interactions where an anti-strange quark ($\bar{s}$) is picked from the nucleon.
Figure 2: Number of incoming neutrinos in SND@LHC as a function of energy. The spectrum is obtained from Ref. Ahdida:2750060 , scaling to an expected integrated luminosity of $\mathcal{L}=290~{}\rm{fb}^{-1}$ and selecting only the neutrinos within the SND@LHC geometrical acceptance.
To this end, the interaction probability of a neutrino with the detector is computed as:
$P_{\rm int}$ $=\displaystyle\frac{{\rm A}\,\sigma_{\nu N}^{\rm model}}{\mathcal{S}}\frac{m_{\rm target}}{m_{\rm N}}\,$ (26)
where $\sigma_{\nu N}^{\rm model}$ is the charge current scattering cross-section for Eqs. (20), (21). The detector consists of 5 walls, each with 2×2 tungsten bricks with an area perpendicular to the beam direction of $\mathcal{S}=400\times 400\,{\rm mm}^{2}$ Ahdida:2750060 . The total mass of the detector is $m_{\rm target}=830$ kg Ahdida:2750060 , and $m_{\rm N}=3.05\times 10^{-25}$ kg is the mass of a tungsten nucleus. The length of the detector is $l_{\rm target}=59\times 5\,$mm, and the mass number of tungsten is A=183.
The cross-section $\sigma_{\nu N}$ is obtained as follows:
The differential cross-section for an incoming neutrino with energy $E_{\nu}$ scattering off a nucleon of mass $M$ in terms of the scattering amplitude is:
$\frac{d\sigma}{dxdy}=\frac{1}{32\pi ME_{\nu}}\int\frac{d\xi}{\xi}f(\xi)|\bar{\mathcal{M}}(\xi)|^{2}\delta(\xi-x).$ (27)
where $\xi$ is the momentum fraction defined by $p_{q}^{\mu}=\xi p^{\mu}$, with $p_{q}^{\mu}$ being the four-momentum of the scattered quark and $p^{\mu}$ the target nucleon momentum. The function $f(\xi)$ represents the parton distribution function (PDF) within the nucleon. We further decompose the differential cross-section of the neutrino-nucleon DIS into contributions from the SM, NP operators, and their interference terms:
$\frac{d\sigma_{\nu N}^{\rm SM+NP}}{dxdy}=\frac{d\sigma_{\nu N}^{\rm SM}}{dxdy}+\frac{d\sigma_{\nu N}^{\rm NP}}{dxdy}+\frac{d\sigma_{\nu N}^{\rm Interference}}{dxdy},$ (28)
where $x$ is the Bjorken variable and $y$ is the inelasticity with $q$ being the four-momentum transfer of the leptonic probe and $\nu\,=-p\cdot q=M(E_{\nu}-E_{\ell})$. The complete expressions for these terms are provided below for different interactions.
Scalar interactions: Using Ref. Liu:2015rqa , we have
$\frac{d\sigma^{\rm SM}}{dxdy}$ $=$ $\frac{G_{F}^{2}ME_{\nu}}{\pi}\left(xy^{2}+\frac{ym_{\ell}^{2}}{2ME_{\nu}})F_{1}+(1-y-\frac{Mxy}{2E_{\nu}}-\frac{m_{\ell}^{2}}{4E_{\nu}^{2}})F_{2}\right.$
$\left.+(xy-\frac{xy^{2}}{2}-\frac{ym_{\ell}^{2}}{4ME_{\nu}})F_{3}-\frac{m_{\ell}^{2}}{2ME_{\nu}}F_{5}\right),$
$\frac{d\sigma^{\rm S}}{dxdy}$ $=$ $\frac{G_{F}^{2}ME_{\nu}}{\pi}(\epsilon_{S}^{2}+\epsilon_{P}^{2})y(xy+\frac{m_{\ell}^{2}}{2ME_{\nu}})F_{1},$
$\frac{d\sigma^{\rm SM,S}}{dxdy}$ $=$ $0$ (29)
The functions $F_{i}$ are given as
$F_{1}$ $=$ $\sum_{q,\bar{q}}f_{q,\bar{q}}(\xi,Q^{2})V^{2}_{q,c},$
$F_{2}$ $=$ $2\sum_{q,\bar{q}}\xi f_{q,\bar{q}}(\xi,Q^{2})V^{2}_{q,c},$
$F_{3}$ $=$ $2\sum_{q}f_{q}(\xi,Q^{2})V^{2}_{q,c}-2\sum_{\bar{q}}f_{\bar{q}}(\xi,Q^{2})V^{2}_{\bar{q},\bar{c}},$
$F_{5}$ $=$ $2\sum_{q,\bar{q}}f_{q,\bar{q}}(\xi,Q^{2})V^{2}_{q,c},$ (30)
where $f_{q}$ and $f_{\bar{q}}$ are the parton distribution functions inside a nucleon, $V_{q,q^{\prime}}$ is the CKM matrix element, and $Q^{2}=-q^{2}$.
Vector and axial vector interactions: For vector and axial interactions, we have
$\frac{d\sigma_{\nu N}^{\rm SM+VA}}{dxdy}\,$ $=$ $\,\frac{G_{F}^{2}ME_{\nu}}{\pi}\left[\left(\frac{|a^{\prime}|^{2}+|b^{\prime}|^{2}}{2}\right)\left(\frac{ym_{\ell}^{2}}{2ME_{\nu}}\,+\,xy^{2}\right)F_{1}(x)\right.$
$\left.+\left(\frac{|a^{\prime}|^{2}+|b^{\prime}|}{2}^{2}\right)\left(1-y-\frac{xyM}{E_{\nu}}-\frac{m_{\ell}^{2}}{4E_{\nu}^{2}}\right)F_{2}(x)\right.$
$\left.+{\rm Re}(a^{\prime}b^{{}^{\prime}*})\left(xy-\frac{xy^{2}}{2}-\frac{ym_{l}^{2}}{4ME_{\nu}}\right)F_{3}(x)-\left(\frac{|a^{\prime}|^{2}+|b^{\prime}|^{2}}{2}\right)\frac{m_{l}^{2}}{2ME_{\nu}}F_{5}(x)\right]$
[…]
The total cross-section $\sigma_{\nu N}$ is obtained by integrating the differential cross-section over the following limits:
$\frac{m_{l}^{2}}{2M(E_{\nu}-m_{l})}\,\leq\,x\,$$\leq$  $\,1,$ (32)
$A-B\,\leq\,y\,\leq\,A+B$ (33)
where
$A\,$ $=$ $\,\frac{1}{2}\left(1-\frac{m_{l}^{2}}{2ME_{\nu}x}-\frac{m_{l}^{2}}{2E_{\nu}^{2}}\right)\Big{/}\left(1+\frac{xM}{2E_{\nu}}\right),$ (34)
$B\,$ $=$ $\,\frac{1}{2}\left[\left(1-\frac{m_{l}^{2}}{2ME_{\nu}x}\right)^{2}-\frac{m_{l}^{2}}{E_{\nu}^{2}}\right]\Big{/}\left(1+\frac{xM}{2E_{\nu}}\right).$ (35)
5 Results
[…]
In principle, NP affecting charm quarks, as encoded in the effective Hamiltonian of Eq. (4), will also impact the production flux of neutrinos. Ref. Kling:2021gos demonstrates that charm-hadron decays predominantly drive the electron-neutrino flux at energies above 200 GeV, while they contribute equally to the muon-neutrino flux, along with pion decays, at energies above 400 GeV. These effects have been thoroughly analyzed in Ref. Falkowski:2021bkq for all neutrino flavors. Their study shows that an NP pseudoscalar interaction with order-one coupling significantly amplifies the flux at production due to 2-body $D_{s}$ decays, increasing the electron-neutrino flux by up to $10^{3}\times$ and the muon-neutrino flux by up to 27×27$\times$27 × than in the SM. However, $\nu_{e}$ and $\nu_{\mu}$ production are dominated by three-body $D_{s}$ decays, which unlike two-body $D_{s}$ decays are not similarly enhanced. Therefore, these production and detection enhancements do not impact our analysis, as we only consider NP in the charm sector. We emphasize that, even if NP were present in first-generation quarks, stringent low-energy constraints impose strong limits on the corresponding Wilson coefficients, rendering current neutrino detectors ineffective in probing such NP operators. Furthermore, in our analysis even at a high luminosity, significant sensitivity to NP affecting muons is only observed when the final state is charm tagged. We, therefore, assume a 100% charm-tagging efficiency.
[…]
Figure 3: Number of signal (SM + NP, dashed) and background (SM only, solid) events as functions of the incoming neutrino energy ($E_{\nu}$) for $\nu_{e}+N\to e^{-}+X_{c}$ (left plot) and $\nu_{\mu}+N\to\mu+X_{c}$ (right plot.) The integrated luminosity ${\cal L}$ has been set to the SND@LHC current projection of 290 fb-1 for the lower (dark blue/dark pink) histograms, and to its high-lumi upgrade projection of 3000 fb-1 for the upper (light blue/light pink) histograms. These results use couplings of model 17 from Tables 2 and 3, alongwith the pseudorapidity range $8.47.2\,<\eta\,<$ 8.4 corresponding to the SND@LHC detector.
The resulting event distributions for the electron and muon neutrinos are presented in Fig. 3. As can be seen in Fig. 3, the NP contribution is negligible for the electron neutrino. In Table 4, we present our results for the sensitivities of detecting NP with muons in the final state. This sensitivity is influenced by the left- and right-handed vector couplings. These results are based on the central values of NP couplings listed in Table 3. Here, we only include those cases where the sensitivity exceeds $0.1\,\sigma$.
Models Sens($\nu_{\mu}+N\to\mu+X_{c}$) Sens$\left(\displaystyle\frac{\nu_{\mu}+N\to\mu+X_{c}}{\nu_{e^{-}}+N\to e^{-}+X_{c}}\right)$
[…]
The high-luminosity upgrade of SND@LHC can collect ten times more data Abbaneo:2895224 , significantly enhancing the sensitivity to NP, even in models where no significant sensitivity is observed at = $\mathcal{L}\,=\,290\,{\rm fb}^{-1}$. We assess the sensitivity for a luminosity of 3000 fb-1 in Table 5. Fig. 3 also includes event distributions for the high-lumi upgrade of SND@LHC.
Models Sens($\nu_{\mu}+N\to\mu+X_{c}$) Sens$\left(\displaystyle\frac{\nu_{\mu}+N\to\mu+X_{c}}{\nu_{e^{-}}+N\to e^{-}+X_{c}}\right)$
[…]
Table 5: Sensitivity of detecting NP in the signal $\nu_{\mu}+N\to\mu+X_{c}$ at SND@LHC, based on models listed in Table 3, assuming a luminosity of 3000 fb-1.
Finally, we note the importance of $\nu_{\tau}$ scattering at experiments such as SND@LHC. NP effective operators involving both the charm quark and the $\tau$ neutrino are difficult to constrain from low-energy measurements. Due to the relative closeness in the $\tau$ and charm masses, three-body decays are suppressed or prohibited due to phase-space effects. In these cases, neutrino scattering experiments can shed a significant amount of light on such NP operators. In this article, however, we restrict ourselves to the cases where neutrino scattering experiments can provide information complementary to low-energy constraints, and as such do not consider $\tau$-neutrino scattering.
6 Conclusion
In this paper, we have presented a detailed sensitivity analysis of new physics (NP) and lepton universality violation (LUV) in neutrino scattering at SND@LHC, focusing on processes involving a charm quark in the final state. Within the framework of effective field theory, NP effects were described using higher-dimensional four-Fermi operators. The Wilson coefficients for these dimension-6 operators were determined from fits to low-energy data, particularly from decays of charmed mesons and baryons. Using benchmark NP points derived from these low-energy analyses, we provide sensitivity estimates for detecting NP and LUV effects at SND@LHC. Both production and scattering processes of neutrinos were considered in evaluating NP effects.
[…]
We also believe that a more comprehensive, multi-dimensional analysis could further improve the sensitivity. Additionally, our findings indicate that these detectors may be sensitive to NP operators that evade low-energy constraints but become relevant at high-energy scales. One such operator, as proposed in Ref. Datta:1996gg , fits this profile, though we leave a detailed investigation for future work. Furthermore, since effective couplings involving charmed particle decays with a final state $\nu_{\tau}$ receive less stringent constraints from low-energy experiments due to phase space restrictions, we note that neutrino scattering experiments can provide useful information about NP in such effective operators.
Acknowledgments: This work was financially supported by the U.S. National Science Foundation under Grant No. PHY-2310627 (BB) and PHY-2309937 (AD). BB acknowledges support in part from the NSF Grant No. PHY-2309135 to the Kavli Institute of Theoretical Physics (KITP) where part of this work was completed. DS is supported by funding from the European Union’s Horizon 2020 research and innovation programme under grant agreement No. 101002846 (ERC CoG “CosmoChart”). The work of EG is supported by the Swiss National Science Foundation (SNSF) under grant number 202065 (“Ambizione”). We are grateful to the SND@LHC collaboration for several useful discussions. We thank J. Kopp, F. Kling, and Z. Tabrizi for pointing out that three-body decays play a vital role in the production of electron and muon neutrinos for SND@LHC.
Appendix A Lepton Universality Violation in charged-current $B$ decays
[…]
Observable SM Prediction Measurement
$R_{D^{*}}^{\tau/\ell}$ $0.254\pm 0.005$ HFLAV:RDRDst2024update $0.287\pm 0.012$ HFLAV:RDRDst2024update
$R_{D}^{\tau/\ell}$ $0.298\pm 0.004$ HFLAV:RDRDst2024update $0.342\pm 0.026$ HFLAV:RDRDst2024update
$R_{J/\psi}^{\tau/\mu}$ $0.283\pm 0.048$ Watanabe:2017mip $0.71\pm 0.17\pm 0.18$ Aaij:2017tyk
$R_{D^{*}}^{\mu/e}$ $\sim 1.0$ $1.04\pm 0.05\pm 0.01$ Belle:2017rcc
Table 6: Standard Model predictions and measured values of observables that may provide hints of LUV NP in charged-current $B$ decays.
Appendix B Hadronic amplitudes for three body decays
The hadronic matrix element of the vector current for the three body decay $M_{i}\to M_{j}\ell\nu$ in terms of the form factors $f_{+,0}(q^{2})$ is given as
$\langle M_{j}(k^{\prime})|\bar{q_{j}}\gamma_{\mu}c|M_{i}(k)\rangle=f_{+}(q^{2})\left((k+k^{\prime})_{\mu}-\frac{m_{M_{i}}^{2}-m_{M_{j}}^{2}}{q^{2}}q_{\mu}\right)+f_{0}(q^{2})\frac{m_{M_{i}}^{2}-m_{M_{j}}^{2}}{q^{2}}q_{\mu}.$ (36)
Using the equation of motion, $i\partial_{\mu}(\bar{q}_{j}\gamma^{\mu}c)=(m_{c}-m_{j})(\bar{q}_{j}c)$, the scalar matrix element can be written as
$\langle M_{j}|\bar{q_{j}}c|M_{i}\rangle=f_{0}(q^{2})\frac{m_{M_{i}}^{2}-m_{M_{j}}^{2}}{m_{c}-m_{j}}.$ (37)
The non-vanishing helicity amplitudes in terms of the form factors are given by
$h_{V,0}^{2}(q^{2})$ $=$ $\sqrt{\frac{\lambda(m_{M_{i}}^{2},m_{M_{j}}^{2},q^{2})}{q^{2}}}f_{+}(q^{2}),$ (38)
$h_{V,t}^{2}(q^{2})$ $=$ $\frac{m_{M_{i}}^{2}-m_{M_{j}}^{2}}{\sqrt{q^{2}}}f_{0}(q^{2}),$ (39)
$h_{S}^{2}(q^{2})$ $\simeq$ $\frac{m_{M_{i}}^{2}-m_{M_{j}}^{2}}{m_{c}-m_{j}}f_{0}(q^{2}).$ (40)
Similarly, the vector and axial vector hadronic matrix elements for the $M_{i}\to M_{j}^{*}\ell\nu$ decays as a function of the form factors can be written as
$\langle M^{*}(k,\varepsilon)|\bar{q}_{j}\gamma_{\mu}c|\overline{M}_{i}(p)\rangle$ $=$ $-i\epsilon_{\mu\nu\rho\sigma}\varepsilon^{\nu*}p^{\rho}k^{\sigma}\frac{2V(q^{2})}{m_{M_{i}}+m_{M_{j}^{*}}},$ (41)
$\langle M^{*}(k,\varepsilon)|\bar{q}_{j}\gamma_{\mu}\gamma_{5}c|\overline{M}_{i}(p)\rangle$ $=$ $\varepsilon^{\mu*}(m_{M_{i}}+m_{M_{j^{*}}})A_{1}(q^{2})-(p+k)_{\mu}(\varepsilon^{*}q)\frac{A_{2}(q^{2})}{m_{M_{i}}+m_{M_{j^{*}}}}$ (42)
$~{}~{}-q_{\mu}(\varepsilon^{*}q)\frac{2m_{M_{j}^{*}}}{q^{2}}(A_{3}(q^{2})-A_{0}(q^{2})),$
where
$A_{3}(q^{2})=\frac{m_{M_{i}}+m_{M_{j^{*}}}}{2m_{M_{j}^{*}}}A_{1}(q^{2})-\frac{m_{M_{i}}-m_{M_{j^{*}}}}{2m_{M_{j}^{*}}}A_{2}(q^{2}).$ (43)
The kinematic constraint at zero recoil gives $A_{3}(0)=A_{0}(0)$. The pseudoscalar matrix element is determined using the equation of motion $i\partial_{\mu}(\bar{q}_{j}\gamma^{\mu}\gamma^{5}c)=-(m_{c}+m_{j})(\bar{q}_{j}\gamma^{5}c)$ and is given by
$\langle M^{*}(k,\varepsilon)|\bar{q}_{j}\gamma_{5}c|\overline{M}_{i}(p)\rangle=-(\varepsilon^{*}q)\frac{2m_{M_{j^{*}}}}{m_{c}+m_{j}}A_{0}(q^{2}).$ (44)
The helicity amplitudes in terms of the form factors are given as
$H_{V,\pm}(q^{2})$ $=$ $(m_{M_{i}}+m_{M_{j^{*}}})A_{1}(q^{2})\mp\frac{\sqrt{\lambda(m_{M_{i}}^{2},m_{M_{j}^{*}}^{2},q^{2})}}{m_{M_{i}}+m_{M_{j}^{*}}}V(q^{2}),$ (45)
$H_{V,0}(q^{2})$ $=$ $\frac{m_{M_{i}}+m_{M_{j^{*}}}}{2m_{M_{j}^{*}}\sqrt{q^{2}}}\bigg{\{}-(m_{M_{i}}^{2}-m_{M_{j}^{*}}^{2}-q^{2})A_{1}(q^{2})$ (46)
$\hskip 85.35826pt+\frac{\lambda(m_{M_{i}}^{2},m_{M_{j}^{*}}^{2},q^{2})}{(m_{M_{i}}+m_{M_{j}^{*}})^{2}}A_{2}(q^{2})\bigg{\}},$
$H_{V,t}(q^{2})$ $=$ $-\sqrt{\frac{\lambda(m_{M_{i}}^{2},m_{M_{j}^{*}}^{2},q^{2})}{q^{2}}}A_{0}(q^{2}),$ (47)
$H_{S}(q^{2})$ $\simeq$ $-\frac{\sqrt{\lambda(m_{M_{i}}^{2},m_{M_{j}^{*}}^{2},q^{2})}}{m_{c}+m_{j}}A_{0}(q^{2}).$ (48)
As mentioned earlier, we incorporate the hadronic matrix elements for the baryonic $\Lambda_{c}\to\Lambda\ell\nu$ decays from Ref. Datta:2017aue .
Appendix C Numerical inputs and form factors
Here we collect the relevant numerical inputs (in Table 7) and form factors for the charmed hadron decays to pions, kaons, and other light pseudoscalar and vector resonances used in the analysis.
$G_{F}$ $1.166\times 10^{-5}$ ParticleDataGroup:2022pth $m_{K^{+}}$ $493.677$ MeV ParticleDataGroup:2022pth
$f_{D_{s}}$ $249.9(5)$ MeV  FlavourLatticeAveragingGroupFLAG:2021npn $m_{K^{0}}$ $497.611$ MeV ParticleDataGroup:2022pth
$f_{D}$ $212.0(7)$ MeV FlavourLatticeAveragingGroupFLAG:2021npn $m_{\pi^{+}}$ $139.57$ MeV ParticleDataGroup:2022pth
$|V_{cs}|$ $0.983(2)(18)$ FlavourLatticeAveragingGroupFLAG:2021npn $m_{\pi^{0}}$ $134.98$ MeV ParticleDataGroup:2022pth
$|V_{cd}|$ $0.2179(7)(57)$ FlavourLatticeAveragingGroupFLAG:2021npn $m_{K^{*+}}$ $891.67$ MeV ParticleDataGroup:2022pth
$\tau_{D_{s}}$ $0.501$ ps ParticleDataGroup:2022pth $m_{K^{*0}}$ $895.55$ MeV ParticleDataGroup:2022pth
$\tau_{D^{+}}$ $1.033$ ps ParticleDataGroup:2022pth $m_{\rho^{+}}$ $766.5$ MeV ParticleDataGroup:2022pth
$\tau_{D^{0}}$ $0.410$ ps ParticleDataGroup:2022pth $m_{\rho^{0}}$ $775.26$ MeV ParticleDataGroup:2022pth
$m_{D_{s}^{+}}$ $1968.35$ MeV FlavourLatticeAveragingGroupFLAG:2021npn $m_{\phi}$ $1.019$ GeV ParticleDataGroup:2022pth
$m_{D^{+}}$ $1869.66$ MeV ParticleDataGroup:2022pth $m_{\eta}$ $547.86$ MeV ParticleDataGroup:2022pth
$m_{D^{0}}$ $1864.84$ MeV ParticleDataGroup:2022pth $m_{\eta^{\prime}}$ $957.78$ Me ParticleDataGroup:2022pth
$m_{D^{*}}$ $2.1122$ GeV ParticleDataGroup:2022pth $m_{\Lambda_{c}}$ $2.286$ GeV ParticleDataGroup:2022pth
$m_{D^{*0}}$ $2.317$ GeV ParticleDataGroup:2022pth $m_{\Lambda}$ $1.116$ GeV ParticleDataGroup:2022pth
Table 7: List of all numerical inputs used in the analysis.
For the $D\to\pi\ell\nu_{\ell}$ transition, the $q^{2}$ dependence of two relevant form factors $f_{+,0}^{D\to\pi}$ can be expressed as Lubicz:2017syv :
$f_{+}^{D\to\pi}(q^{2})$ $=$ $\frac{1}{1-P_{V}q^{2}}\bigg{\{}f^{D\to\pi}(0)+c_{+}^{D\to\pi}(z-z_{0})\left(1+\frac{z+z_{0}}{2}\right)\bigg{\}},\,$ (49)
$f_{0}^{D\to\pi}(q^{2})$ $=$ $\frac{1}{1-P_{S}q^{2}}\bigg{\{}f^{D\to\pi}(0)+c_{0}^{D\to\pi}(z-z_{0})\left(1+\frac{z+z_{0}}{2}\right)\bigg{\}},\,$ (50)
where $z$ is defined as
$z=\frac{\sqrt{t_{+}-q^{2}}-\sqrt{t_{+}-t_{0}}}{\sqrt{t_{+}-q^{2}}+\sqrt{t_{+}-t_{0}}},$ (51)
with
$t_{+}$ $=$ $(m_{D}+m_{\pi})^{2},$ (52)
$t_{0}$ $=$ $(m_{D}+m_{\pi})(\sqrt{m_{D}}-\sqrt{m_{\pi}})^{2}.$ (53)
The other parameters that appear in the z-expansion of the form factors are collected below in Table 8.
$f^{D\to\pi}(0)$ $c_{+}^{D\to\pi}$ $P_{V}~{}({\rm GeV}^{-2})$ $c_{0}^{D\to\pi}$ $P_{S}~{}({\rm GeV}^{-2})$
 0.6117(354)  -1.985(347)  0.1314(127)  -1.188(256)  0.0342(122)
Table 8: Values of the parameters taken from Lubicz:2017syv that appear in the z-expansion of the vector and scalar form factors of the $D\to\pi$ transition.
A similar kind of z-expansion is also employed in case of the $D\to K$ transition are expressed as
$f_{+}^{D\to K}(q^{2})$ $=$ $\frac{1}{1-q^{2}/m_{D^{*}}^{2}}\left(a_{0}^{+}+a_{1}^{+}(z-z^{3}/3)+a_{2}^{+}(z^{2}+2z^{3}/3)\right),$ (54)
$f_{0}^{D\to K}(q^{2})$ $=$ $\frac{1}{1-q^{2}/m_{D^{*0}}^{2}}\left(a_{0}^{0}+a_{1}^{0}~{}z+a_{2}^{0}~{}z^{2})\right),$ (55)
with $m_{\pi}\leftrightarrow m_{K}$ in Eqs. (52)-(53). The kinematic constraint at zero recoil $f_{+}^{D\to K}(0)=f_{0}^{D\to K}(0)$ reduces one parameter out of the six while rest of the coefficients are taken from the most recent FLAG $N_{f}=2+1+1$ computation of the form factors FlavourLatticeAveragingGroupFLAG:2021npn as listed in Table 9.
$a_{0}^{+}$ $a_{1}^{+}$ $a_{2}^{+}$ $a_{0}^{0}$ $a_{1}^{0}$
 0.7877(87)  -0.97(18)  -0.3(2.0)  0.6959(47)  0.775(69)
Table 9: Values of form factor parameters relevant to the $D\to K$ transition as obtained by FlavourLatticeAveragingGroupFLAG:2021npn .
The form factors for the $D_{s}\to(\eta,\eta^{\prime})\ell\nu$ transition are calculated based on the light cone QCD sum rules (LCSR) as in Ref. Azizi:2010zj . The $q^{2}$ dependence of the vector form factors is given by the form
$f_{\pm}^{D_{s}\to\eta}(q^{2})=\frac{f_{\pm}(0)^{D_{s}\to\eta}}{1-\alpha_{\pm}\hat{q}+\beta_{\pm}\hat{q}^{2}},$ (56)
with $\hat{q}=q^{2}/m_{D_{s}}^{2}$ and the parameters $\alpha_{\pm}$, $\beta_{\pm}$ and $f_{\pm}(0)^{D_{s}\to\eta}$ are listed in Table. 10. The scalar form factor $f_{0}^{D_{s}\to\eta}$ is related to the vector ones by the relation
$f_{0}^{D_{s}\to\eta}(q^{2})=f_{+}^{D_{s}\to\eta}(q^{2})+\frac{q^{2}}{m_{D_{s}}^{2}-m_{\eta}^{2}}f_{-}^{D_{s}\to\eta}(q^{2}).$ (57)
The form factors $f_{i}^{D_{s}\to\eta^{\prime}}$ are related to $f_{i}^{D_{s}\to\eta}$ as Azizi:2010zj
$\frac{|f_{i}^{D_{s}\to\eta}|}{|f_{i}^{D_{s}\to\eta^{\prime}}|}=\tan\varphi,$ (58)
where $\varphi=(39.7\pm 0.7)^{\circ}$ Azizi:2010zj is the mixing angle between the $\eta$ and $\eta^{\prime}$ states.
$f_{+}^{D_{s}\to\eta}(0)$ $\alpha_{+}$ $\beta_{+}$ $f_{-}^{D_{s}\to\eta}(0)$ $\alpha_{-}$ $\beta_{-}$
 0.45(14)  1.96(63)  1.12(36)  -0.44(13)  2.05(65)  1.08(35)
Table 10: Values of form factor parameters relevant to the $D_{s}\to\eta$ transition as obtained by the LCSR calculation in Ref. Azizi:2010zj .
A simple single pole parametrization is employed to calculate the form factors for the pseudoscalar $P=D,D_{s}$ to vector $M=K^{*},\phi,\rho$ meson transitions given as
$F_{i}^{M}(q^{2})=\frac{F_{i}^{M}(0)}{1-q^{2}/m_{R,i}^{2}},$ (59)
where $F_{i}=A_{1},A_{2},A_{0},V$ and $m_{R,i}$ are the respective pole masses. The kinematic constraint $A_{0}(0)=A_{3}(0)=(\frac{m_{P}+m_{M}}{2m_{M}}A_{1}(0)-\frac{m_{P}-m_{M}}{2m_{M}}A_{2}(0))$ relates the parameters at zero recoil. We list the values of the relevant parameters in Table. 11.
$F_{i}$ $K^{*}$ BaBar:2010vmf $\rho$ Fu:2018yin $\phi$ Donald:2013pea
$F_{i}(0)$ $m_{R,i}$ (GeV) $F_{i}(0)$ $m_{R,i}$ (GeV) $F_{i}(0)$ $m_{R,i}$ (GeV)
 $A_{1}$  0.620(85)  2.459  $0.58^{+0.065}_{-0.050}$  2.427  0.615(24)  2.459
 $A_{2}$  $r_{2}A_{1}(0)$ 2.459 $0.468^{+0.052}_{-0.053}$ 2.427 0.457(78) 2.459
$r_{2}=0.801(30)$
 $V$  $r_{V}A_{1}(0)$ 2.112 $0.815^{+0.070}_{-0.051}$ 2.007 1.059(124) 2.112
$r_{V}=1.463(35)$
Table 11: Values of form factor parameters used in our analysis for the $P\to M$ transitions along with the reference from which they are taken.
For the semileptonic baryonic decay, we employ the most recent lattice QCD computation of the form factors reported in Meinel:2016dqj . There are a total of six form factors $f_{+},f_{\perp},f_{0},g_{+},g_{\perp},g_{0}$ which are parameterized using a simple pole z-expansion of the form
$f(q^{2})=\frac{1}{1-q^{2}/(m_{\rm pole}^{f})^{2}}\sum_{n=0}^{n_{\rm max}}a_{n}^{f}\left[z(q^{2})\right]^{n},$ (60)
where z is defined as in Eq. (51) with $t_{0}=q^{2}_{\rm max}=(m_{\Lambda_{c}}-m_{\Lambda})^{2}$ and $t_{+}=(m_{D}+m_{K})^{2}$. The pole masses are $m_{\rm pole}^{f_{+},f_{\perp}}=2.112$ GeV, $m_{\rm pole}^{f_{0}}=2.318$ GeV, $m_{\rm pole}^{g_{+},g_{\perp}}=2.460$ GeV and $m_{\rm pole}^{g_{0}}=1.968$ GeV. The “nominal fit” results for the form factor parameters provided in  Meinel:2016dqj have been used for our analysis.